# SpecCLIP: Aligning and Translating Spectroscopic Measurements for Stars

### v1

@@ 6 Summary
  • Our models support both in-modality and cross-modality spectrum retrieval, as well as spectrum-to-spectrum prediction across modalities. High similarity and prediction scores demonstrate that the learned representations capture shared physical information. These modules also offer promising avenues for anomaly detection and similarity-based searches in large spectral archives.
@@ Acknowledgements
Here we present additional information on multiple aspects of this work, including elemental-abundance prediction, continuum fitting, normalization flows for parameter estimation, the use of pre-trained models, and projection models and decoders.
@@
Table 5: Similar to Table 1 for “LRS Models”, but the results for all CLIP-based models are from the alignment between the LAMOST LRS MT and Gaia XP MT (see Section 4.1) for all model variants.
@@
$\mathrm{[\alpha/Fe]}$ 0.023 / --2.376 0.022 / 0.901 0.021 / 0.908 0.020 / 0.909 0.021 / 0.908 0.021 / 0.906
$\mathrm{[C/Fe]}$ 0.039 / 0.000 0.039 / 0.784 0.038 / 0.798 0.038 / 0.799 0.038 / 0.801 0.038 / 0.795
$\mathrm{[N/Fe]}$ 0.054 / --6.346 0.052 / 0.630 0.051 / 0.653 0.051 / 0.654 0.051 / 0.659 0.051 / 0.657
$\mathrm{[Al/Fe]}$ 0.048 / --2.404 0.049 / 0.703 0.046 / 0.743 0.047 / 0.739 0.047 / 0.743 0.047 / 0.737
$\mathrm{[Ca/Fe]}$ 0.031 / --8.202 0.031 / 0.691 0.030 / 0.705 0.030 / 0.707 0.030 / 0.707 0.030 / 0.707
$\mathrm{[Mg/Fe]}$ 0.031 / --4.301 0.032 / 0.867 0.031 / 0.876 0.031 / 0.877 0.031 / 0.879 0.032 / 0.874
$\mathrm{[Si/Fe]}$ 0.030 / --5.773 0.029 / 0.798 0.028 / 0.806 0.028 / 0.807 0.029 / 0.808 0.029 / 0.802
$\mathrm{[Ti/Fe]}$ 0.058 / --0.339 0.058 / 0.528 0.057 / 0.547 0.057 / 0.547 0.057 / 0.547 0.057 / 0.537
$\mathrm{[Mn/Fe]}$ 0.034 / --6.203 0.032 / 0.779 0.031 / 0.796 0.031 / 0.797 0.031 / 0.798 0.031 / 0.792
$\mathrm{[Ni/Fe]}$ 0.026 / --0.083 0.026 / 0.450 0.025 / 0.483 0.025 / 0.482 0.025 / 0.486 0.026 / 0.476
$\mathrm{[O/Fe]}$ 0.052 / --4.127 0.051 / 0.717 0.050 / 0.723 0.050 / 0.724 0.050 / 0.723 0.050 / 0.722
$\mathrm{[Cr/Fe]}$ 0.078 / --0.575 0.076 / 0.222 0.075 / 0.233 0.075 / 0.234 0.075 / 0.234 0.075 / 0.230
@@
$E(BP-RP)$ 0.076 / --23.199 0.077 / 0.708 0.071 / 0.743 0.070 / 0.743 0.070 / 0.745 0.070 / 0.736
$v_{r}$ (km s-1) 6.418 / 0.942 5.945 / 0.977 6.736 / 0.970 6.572 / 0.970 6.159 / 0.973 5.415 / 0.979
$v_{r}$-sbi(maf) (km s-1) 4.715 / 0.967 4.692 / 0.978 5.815 / 0.956 5.691 / 0.959 5.022 / 0.971 4.687 / 0.979
@@ Appendix A Additional Results for Elemental-abundance Estimation
Figure 6 presents an example of chemical-abundance predictions from one input spectrum. This result is obtained using SBI trained with a single model for all parameters simultaneously. The inset in the upper right corner shows results from simulation-based calibration (SBC). Most elemental abundances exhibit well-calibrated posteriors, except for [Ti/Fe] and [Ca/Fe], which fall below the 0.05 significance threshold. This suggests that, although the model learns posterior distributions well overall, further tuning of the SBI architecture or its hyperparameters is necessary to achieve reliable multivariate inference of elemental abundances.
In Table 5, we provide downstream task results using CLIP-based models aligned between the LAMOST LRS MT and Gaia XP MT encoders. This serves as a comparison to Table 1, where the alignment is performed between the LAMOST LRS MT and Gaia XP OAE models. We find that the performance differences between these two configurations are negligible across the evaluation metrics used.
@@ Appendix B Continuum Fitting Algorithm
    Denoising with Savitzky–Golay filter: The Savitzky–Golay filter is applied to each segment independently, with a smoothing window size adaptively chosen based on the input SNR.
@@ D.1.1 Masked Spectral Modeling for LAMOST Spectra
LAMOST spectra benefit from local patterns and detailed features. Thus, using dense chunk-based masking and slicing helps the transformer leverage locality while preserving the global context. We train this model with 8 A100 GPUs over a total of 128 epochs, a process that requires roughly 20 hours.
@@ D.1.2 Masked Spectral Modeling for Gaia XP Spectra
We train this model with 8 A100 GPUs and a total of 191 epochs, requiring roughly 40 hours.
@@ Appendix E Projection Models and Decoders
This section provides details about the modules used in CLIP-based model training. For these models, we use 8 A100 GPU and require a total of roughly 3 hours for training. All model variants use a batch size of 1024 per GPU for contrastive training with decoders. We adopt the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.05. The learning rate is scheduled using a cosine annealing schedule with linear warm-up.
@@ E.2 Decoders
For the CrossModalDecoder of the CLIP-split model, we takes only the shared representation as input. The decoder architecture depends on the output dimension:

### v2

@@ 6 Summary
  • Our models support both in-modal and cross-modal spectrum retrieval, as well as spectrum-to-spectrum prediction across modalities. High similarity and prediction scores demonstrate that the learned representations capture shared physical information. These modules also offer promising avenues for anomaly detection and similarity-based searches in large spectral archives.
@@ Acknowledgements
G.X. and X.T. acknowledge the support from Key R&D Program of Zhejiang (2024SSYS0006).
@@
Here we present additional information on multiple aspects of this work, including elemental-abundance prediction, continuum fitting, normalization flows for parameter estimation, the use of pre-trained models, projection models and decoders, and loss curves.
@@
Table 5: Similar to Table 1 for “LRS Models”, but the results for all CLIP-based models are from the alignment between the LAMOST LRS MT and Gaia XP MT (see Section 5.6 for discussion, and Appendices D and E for model details).
@@
$\mathrm{[\alpha/Fe]}$ 0.023 / 0.872 0.021 / 0.906 0.020 / 0.912 0.020 / 0.912 0.020 / 0.913 0.021 / 0.909
$\mathrm{[C/Fe]}$ 0.041 / 0.758 0.039 / 0.792 0.038 / 0.806 0.038 / 0.804 0.038 / 0.805 0.038 / 0.803
$\mathrm{[N/Fe]}$ 0.054 / 0.598 0.052 / 0.642 0.050 / 0.662 0.050 / 0.661 0.050 / 0.665 0.051 / 0.657
$\mathrm{[Al/Fe]}$ 0.049 / 0.691 0.048 / 0.711 0.046 / 0.739 0.046 / 0.739 0.046 / 0.740 0.046 / 0.737
$\mathrm{[Ca/Fe]}$ 0.032 / 0.670 0.030 / 0.697 0.029 / 0.714 0.029 / 0.714 0.029 / 0.716 0.030 / 0.711
$\mathrm{[Mg/Fe]}$ 0.031 / 0.866 0.032 / 0.871 0.031 / 0.882 0.031 / 0.882 0.031 / 0.881 0.031 / 0.878
$\mathrm{[Si/Fe]}$ 0.029 / 0.776 0.029 / 0.803 0.028 / 0.814 0.028 / 0.813 0.028 / 0.816 0.028 / 0.807
$\mathrm{[Ti/Fe]}$ 0.061 / 0.492 0.058 / 0.532 0.056 / 0.551 0.056 / 0.552 0.056 / 0.555 0.056 / 0.544
$\mathrm{[Mn/Fe]}$ 0.033 / 0.761 0.032 / 0.780 0.031 / 0.800 0.031 / 0.799 0.031 / 0.798 0.031 / 0.792
$\mathrm{[Ni/Fe]}$ 0.027 / 0.426 0.026 / 0.454 0.025 / 0.487 0.025 / 0.489 0.025 / 0.489 0.026 / 0.479
$\mathrm{[O/Fe]}$ 0.051 / 0.698 0.050 / 0.722 0.049 / 0.728 0.049 / 0.729 0.049 / 0.729 0.049 / 0.728
$\mathrm{[Cr/Fe]}$ 0.081 / 0.177 0.076 / 0.225 0.075 / 0.234 0.075 / 0.233 0.075 / 0.237 0.075 / 0.232
@@
$E(BP-RP)$ 0.075 / --36.886 0.076 / 0.711 0.070 / 0.746 0.069 / 0.748 0.069 / 0.748 0.070 / 0.742
$v_{r}$ (km s-1) 6.071 / 0.970 5.345 / 0.978 6.785 / 0.969 6.834 / 0.969 6.226 / 0.973 5.361 / 0.979
$v_{r}$-sbi(maf) (km s-1) 4.573 / 0.963 4.653 / 0.979 5.636 / 0.958 5.621 / 0.963 5.070 / 0.965 4.578 / 0.981
Note. Numbers in bold indicate the best performance (i.e., lowest $\sigma$ or highest $R^{2}$) for each parameter across all models. Results marked exclude a failed NSF sampling case on one extreme spectrum.
Figure 7: Similar to Figure 2, but the results for all CLIP-based models are from the alignment between the LAMOST LRS MT and Gaia XP MT (see Section 5.6 for discussion, and Appendices D and E for model details).
@@ Appendix A Additional Results for Elemental-abundance Estimation
Figure 6 presents an example of chemical-abundance predictions from one input spectrum. This result is obtained using SBI trained with a single model for all parameters simultaneously. The inset in the upper right corner shows results from simulation-based calibration (SBC). Most elemental abundances exhibit well-calibrated posteriors, except for [Ti/Fe] and [Ca/Fe], which fall below the 0.05 significance threshold. This provides an example where, despite the model learns posterior distributions well overall, further tuning of the SBI architecture or its hyperparameters is necessary to achieve reliable multivariate inference of elemental abundances–an aspect we did not further refine in this work.
In Table 5, we provide downstream task results using CLIP-based models aligned between the LAMOST LRS MT and Gaia XP MT encoders. This serves as a comparison to Table 1, where the alignment is performed between the LAMOST LRS MT and Gaia XP OAE models. We find that the latter configuration shows overall comparable or better performance across the evaluation metrics, possibly due to the stronger pretraining of the XP OAE model. For completeness, we include in Figure 7 an external comparison analogous to Figure 2 in the main text, but using the models from Table 5.
@@ Appendix B Continuum Fitting Algorithm
    Denoising with Savitzky–Golay filter: The Savitzky–Golay filter is applied to each segment independently, with a smoothing window size of 3.
@@ D.1.1 Masked Spectral Modeling for LAMOST Spectra
LAMOST spectra benefit from local patterns and detailed features. Thus, using dense chunk-based masking and slicing helps the transformer leverage locality while preserving the global context. We train this model with 8 GPUs over a total of 128 epochs, a process that requires roughly 20 hours.
@@ D.1.2 Masked Spectral Modeling for Gaia XP Spectra
We train this model with 8 GPUs and a total of 191 epochs, requiring roughly 40 hours.
@@ Appendix E Projection Models and Decoders
This section provides details about the modules used in CLIP-based model training. For these models, we use 8 GPU and require a total of roughly 3 hours for training. All model variants use a batch size of 1024 per GPU for contrastive training with decoders. We adopt the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.05. The learning rate is scheduled using a cosine annealing schedule with linear warm-up.
@@ E.2 Decoders
For the cross-modal decoder of the CLIP-split model, we takes only the shared representation as input. The decoder architecture depends on the output dimension:
@@
Appendix F Loss Curves
This section presents the loss curve comparisons between the CLIP-pr and CLIP-split models, as shown in Figure 8. While the CLIP-pr model achieves a lower CLIP loss during training, it exhibits lower absolute cosine similarity on test pairs, as shown in Table 2. This discrepancy may arise from two factors. First, the geometry of the high-dimensional embedding spaces—768 dimensions in the CLIP-pr projected embedding versus 512 in the CLIP-split shared embedding—tends to yield lower cosine similarity in the former. Second, the CLIP loss focuses on relative alignment rather than absolute similarity. Additionally, the CLIP-pr model yields slightly lower cross-modal prediction losses. In contrast, the CLIP-split model converges more quickly in learning the reconstruction and achieves marginally lower reconstruction loss, likely due to its design of a non-shared projected embedding space.
Figure 8: Loss curves of the CLIP-pr and CLIP-split models.